A tank fo water at 2.0atm is compressed till it reaches 10.5atm. If the orginal temperature of the water vapor was 115 degrees C what is the new temp? Is it Boyles, Charles or Gay-Lussac's law and solve it.
Answer #1

This will e solved by the Gay-Lussac's law P1/T1= P2/T2 or T2 = P2T1/P1

here P1 = 2 atm P2 = 10.5 atm T1= 115 C+273.15=388.15 K ; T2 = ?

using the values

T2= 10.5*388.15/2= 2037.78 K = 1764.63 degree C